OpenStudy (anonymous):

reduce to lowest terms 15/65

OpenStudy (anonymous):

divide by 5

OpenStudy (anonymous):

What is the common number that you can multiply to get 15 and 65?

OpenStudy (anonymous):

The common thing between those the numbers that they can be divided by 5

OpenStudy (anonymous):

5?

OpenStudy (anonymous):

15/5 =3 65/5=13 3/13

OpenStudy (anonymous):

\[\frac{ 15\div5 }{ 65\div5 }\]

OpenStudy (anonymous):

thanks